OpenStudy (anonymous):

I think I just lost my mind. I can't think of what to do with this problem. I am on the last step: Find the line perpendicular to 4y+3x-2=0 that passes through the points (1,-1) The line perpendicular has a slope of -3/4

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

so... -1= -3/4(1)+b

OpenStudy (anonymous):

and I have to find b but I can't think of what to do. Do I just use algebra?

OpenStudy (anonymous):

Yes. I havent checked your method, but yes you have to use algebra. That will get you the y intercept i believe.

OpenStudy (anonymous):

b=-1/4

OpenStudy (anonymous):

so I think my answer is y=-3/4-1/4